OpenStudy (anonymous):

g(f(h(-3))) f(x)= 2x g(x)= -3x+4 h(x)= 2x-8 can someone please give me the answer?.. ive been stuck for days on this and I really need help..

OpenStudy (amistre64):

what is the value of h(-3) ?

OpenStudy (anonymous):

I don't know... hmm.. that's just the problem all typed out word for word :(

OpenStudy (amistre64):

when we look at h(x) it is defined for values of x when x = -3, what does h equal?

OpenStudy (anonymous):

-3.. right?

OpenStudy (amistre64):

h= 2x-8, when x = -3, id say it is -6-8

OpenStudy (anonymous):

so then what do I do with -6-8? or would that be the answer?

OpenStudy (amistre64):

this is a process, we are asked to find an overall solution, but we have to pick it apart.

OpenStudy (amistre64):

g(f(h(-3))) find h(-3) find f(h(-3)) then we can determine g(f(h(-3)))

OpenStudy (amistre64):

h = 2x - 8, when x=-3, h=-6-8 f = 2x, when x = -6-8, what is f? g = -3x+4, what is g when x=f?

OpenStudy (anonymous):

f=-12x-16x?

OpenStudy (amistre64):

no, -6-8 = -14 f = 2x, when x=-14 f = 2(-14)

OpenStudy (anonymous):

so then -3x+4(-28x) -84x-112x= 196x^2?

OpenStudy (amistre64):

g = -3x+4, what is g when x=f=-28 g = -3(-28) + 4

OpenStudy (anonymous):

so then the answer would be.. -80!?

OpenStudy (amistre64):

84 + 4, your basic math skills need alot of work :)

OpenStudy (anonymous):

hahha I know!! this has been so hard for me.. so the final answer is 88..